The IE of a hydrogen-like bohr atom is 4 Rydberg. What is the radius of the first orbit of the atom?

Grade:11

1 Answers

Vikas TU
7 years ago
From the energy formulae,
|-13.6z^2/n^2| = 4R
13.6z^2 = 4*1.01 * 10^-6
on solving 
z = 2.
Now radius,
r = 0.529 * 1^2/z = 0.529/2 = 0.25 Angstrom or
                                        = 0.25 * 10^-10 meter.